OpenStudy (anonymous):

WILL WIN MEDAL!! What is the area of a circle with a diameter of 16? (Express your answer using π.)

OpenStudy (anonymous):

we know that area of circle = pi r^2 where r is the radius and radius is half of the diameter so radius will be 16/2 = 8 so area will be pi*r^2 =\[64\pi \]

OpenStudy (anonymous):

Thank you! You won yourself a medal!